OpenStudy (anonymous):

Help on : if y = f(x) a function, satisfies the differential equation dy/dx = -4y and f(0) = 6, then f(x) = Am I suppose to solve this using integration? If so, how am I suppose to do it??

OpenStudy (anonymous):

I know that f(0) means that the C would be 6, but how do I solve the antiderivative of that thing?

OpenStudy (anonymous):

\[\frac{ dy }{ y }=-4dx\] integrating \[\int\limits \frac{ 1 }{y }dy=-4 \int\limits dx+c\] \[\ln y=-4x+c\] when x=0,y=6 \[\ln 6=0+c,c=\ln 6\] substitute the value of c and finally y=f(x)

OpenStudy (anonymous):

ah... so this is just like implicit differentiation, only the reverse?

OpenStudy (anonymous):

what I do to the one side I must do for the other to be equal, is that how we can derive the equation?? thanks

OpenStudy (anonymous):

but I think I am suppose to solve for y. Since y is in ln y, how do I find y again?

OpenStudy (anonymous):

\[\ln y=-4x+\ln 6\] \[\ln y-\ln 6=-4x,\ln \frac{ y }{ 6 }=-4x,y=6e ^{-4x}\] \[or~f \left( x \right)=6e ^{-4x}\]

OpenStudy (anonymous):

Ah... Ok, those e rules I must remember... Thank you very much for your help.

OpenStudy (anonymous):

yw